OpenStudy (anonymous):

What is the equation of the line described below written in slope-intercept form? the line passing through point (2, 2) and perpendicular to the line whose equation is y = x

OpenStudy (anonymous):

For perpendicular lines: m1*m2=-1 where m1 and m2 are the slopes of the perpendicular lines

OpenStudy (anonymous):

y+x-4 = 0

OpenStudy (anonymous):

so y=x-4?

OpenStudy (anonymous):

Nope.